OpenStudy (anonymous):

On the coordinate grid of a map. Josie's house is located at (2,7). Her school is located at (-5,5). If each map unit equals one mile, what is the approximate distance from her house to school? Please help!!!!!!

OpenStudy (anonymous):

2^2+7^2, then take sqrt of that.

OpenStudy (anonymous):

so sqrt 54.

OpenStudy (anonymous):

just draw urself a right angled triangle.

OpenStudy (anonymous):

and then use common equation of c^2=a^2+b^2

OpenStudy (anonymous):

about 7.34 miles

OpenStudy (anonymous):

49+4=54?

OpenStudy (anonymous):

LOL! sorry!!!!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yo my bad im helping like 8-9 people?...aha ists 53 ;) common now i kenw that ;)

OpenStudy (anonymous):

lol...its ok. thanks! ;)